# Spike No More: Stabilizing the Pre-training of Large Language Models

### v3

@@ 1 Introduction
In this research, we provide theoretical analyses focusing on the loss spike problem during LLM pre-training. We identify the upper bound of the gradient norms for the Transformer-based LLMs through analyses on the spectral norms of the Jacobian matrices for the sub-layers. If the upper bound is large, the gradients may spike suddenly, and we assume that this phenomenon causes the loss spike. Then, we indicate that the upper bound is large in the typical setting, such as the widely used implementation, Megatron-LM (Shoeybi et al., 2020), and thus, the loss spike is likely to occur. In addition, to make the upper bound sufficiently small, we introduce two conditions: (1) initializing the parameters of sub-layers with a small value and (2) making the standard deviation of each embedding close to $1$. The former condition can be satisfied by the widely used initialization method for LLMs (Shoeybi et al., 2020; Le Scao et al., 2022; Biderman et al., 2023). On the other hand, the latter condition was satisfied in the original Transformer by scaling embeddings (Vaswani et al., 2017), but such scaling is missing from recent implementations. To sum up, through theoretical analyses, we re-evaluate several previous techniques in terms of the stabilization of LLM pre-training.
@@ 2.1 Pre-LN Transformer
where $\mathrm{LN}$ is the layer normalization function [^1] We discuss the difference from the architecture using Root Mean Square layer normalization (RMSNorm) (Zhang & Sennrich, 2019) instead of LN in Appendix D, and the original Transformer architecture, i.e., Post-LN Transformer in Appendix J.. We call the first terms in Equations (1) and (2), i.e., $x$ and $x^{\prime}$, shortcut. In addition, the feed-forward network ($\mathrm{FFN}$) and multi-head self-attention ($\mathrm{Attn}$) are defined as follows [^2] To simplify equations, we omit bias terms.:
@@ 3 Analyses on Gradients of Sub-Layers
Let $x$ and $x^{\prime}$ be the input and intermediate vectors of each layer. Moreover, let $W_{*}$ denote the model parameter matrix in each layer. We assume that $x$, $x^{\prime}$, and $W$ for all layers follow a normal distribution with a mean of 0, i.e., $\mu=0$.
This assumption is valid when we initialize parameters with the normal distribution, the number of heads in Equation (4) is $1$, and $\mathcal{F}$ is an identity function. Empirically, the outputs of each sub-layer are close to the normal distribution as illustrated in Appendix F.
@@ 3.1 Jacobian Matrix of FFN
$\bigg{\|}\frac{\partial y}{\partial x^{\prime}}\bigg{\|}_{2}$ $=\bigg{\|}\frac{\partial(x^{\prime}+\mathrm{FFN}(\mathrm{LN}(x^{\prime})))}{\partial x^{\prime}}\bigg{\|}_{2}=\bigg{\|}I+\frac{\partial(\mathrm{FFN}(\mathrm{LN}(x^{\prime})))}{\partial x^{\prime}}\bigg{\|}_{2}.$ (8)
@@
Regarding the $\mathrm{FFN}$ part, we assume that the activation function $\mathcal{F}$ is an identity function [^3] Appendix G discusses the case where we use the $\mathrm{ReLU}$ as the activation function, which results in the same conclusion with a slightly different coefficient. to simplify the discussion. Under this assumption, the following equation holds:
$\bigg{\|}\frac{\partial\mathrm{FFN}(\mathrm{LN}(x^{\prime}))}{\partial\mathrm{LN}(x^{\prime})}\bigg{\|}_{2}=\|W_{1}W_{2}\|_{2}.$ (10)
Therefore, we can straightforwardly derive the relation $\|W_{1}W_{2}\|_{2}\leq\|W_{1}\|_{2}\|W_{2}\|_{2}$ from the submultiplicativity of the spectral norm. Furthermore, let $\sigma_{1}$ and $\sigma_{2}$ be the standard deviations of $W_{1}$ and $W_{2}$, respectively. From Assumption 1, the spectral norms of $W_{1}$ and $W_{2}$ are obtained by their standard deviations and dimensions (Vershynin, 2018), i.e., $\|W_{1}\|_{2}\approx\sigma_{1}(\sqrt{d}+\sqrt{d_{\mathrm{ffn}}})$ and $\|W_{2}\|_{2}\approx\sigma_{2}(\sqrt{d}+\sqrt{d_{\mathrm{ffn}}})$. Finally, we can express an upper bound of the spectral norms of the Jacobian matrices for $\mathrm{FFN}$ as the following inequality:
@@
where the right-hand side has the relation $\sigma_{1}\sigma_{2}(\sqrt{d}+\sqrt{d_{\mathrm{ffn}}})^{2}=\|W_{1}\|_{2}\|W_{2}\|_{2}$.
@@
We consider the variance ($\mathrm{var}$) of each element in the matrix $zz^{\top}$. Since $z_{i}z_{i}$ follows $\mathcal{X}^{2}$ with $1$ degree of freedom, and $z_{i}z_{j}(i\neq j)$ is the multiplication of two independent values following the standard normal distribution, the variances are as follows:
@@ 3.2 Jacobian Matrix of Self-Attention
Let $Z(\cdot)=\mathrm{concat}(\mathrm{head}_{1}(\cdot),...,\mathrm{head}_{h}(\cdot)))$ and let $J^{Z}$ be the Jacobian of the $Z(\cdot)$ [^4] We discuss the detail of $J^{Z}$ in Appendix I., we can rewrite the spectral norm of the Jacobian matrix of Attn as:
@@ 4 Conditions to Avoid Spikes
However, in this situation, the standard deviation of the shortcut is also too small. For example, at shallow layers, the standard deviation is close to $\sqrt{\frac{2}{5d}}$ because the embedding matrix is also initialized by $\mathcal{N}(0,\sigma^{2})$ where $\sigma=\sqrt{\frac{2}{5d}}$. Therefore, to increase the standard deviation of the shortcut, we make the standard deviation of each embedding close to 1111666Based on Equations (15) and (20), the upper bound becomes small as the standard deviation of the shortcut increases. However, a too large value degrades the performance empirically as described in Appendix E.. To achieve this, we introduce two kinds of modification: “Scaled Embed” and “Embed LN” [^7] We can satisfy the condition by initializing embeddings with the normal distribution $\mathcal{N}(0,\sigma^{2})$ where $\sigma=1$, but we do not adopt this strategy in this study because we use the same initialization method in our experiments.. The Scaled Embed scales embeddings with an appropriate value. For example, we multiply embeddings by $\sqrt{d}$, which was used in the original Transformer paper (Vaswani et al., 2017) [^8] Although the original Transformer paper introduced this operation, recent implementations ignore this., and then the standard deviations of embeddings become $\sqrt{\frac{2}{5}}$. The Embed LN applies the LN to embeddings. In fact, Le Scao et al. (2022) reported that the Embed LN strategy prevents loss spikes empirically. These two methods are presented as verification examples rather than proposed methods, and alternative approaches could be employed if the conditions are met.
@@
To demonstrate the actual values of the upper bound described in Equation (15), we take the model with 1.7 billion parameters as an example. In addition to the widely used initialization for LLMs (Vanilla) and the above two modifications: Scaled Embed and Embed LN, we compare Xavier Init, which initializes all parameters with the Xavier initialization (Glorot & Bengio, 2010), as the situation where we do not scale $W_{2}$ and $W_{O}$ based on the number of layers. Figure 2 shows the values of Equation (15) for each layer at the beginning of the pre-training. This figure indicates that the methods without suppressing the upper bound, i.e., Xavier Init and Vanilla, rapidly increase the values especially in shallow layers. In contrast, Scaled Embed and Embed LN keep small values. In summary, to make the upper bound of the gradient norms small for the stabilization of the LLM pre-training, we have to satisfy two conditions: (1) small sub-layers; initializing the parameters of sub-layers with a small value and (2) large shortcut; making the standard deviation of each embedding close to $1$.
Table 1: Relations between each method in experiments and two conditions to control the upper bound of gradient norms. We conduct experiments for Xavier Init and Xavier Init + Scaled Embed in Appendix B.
Method Small sub-layers Large shortcut
Xavier Init - -
Vanilla $\checkmark$ -
Embed Detach $\checkmark$ -
Embed LN $\checkmark$ $\checkmark$
Scaled Embed $\checkmark$ $\checkmark$
Xavier Init + Scaled Embed - $\checkmark$
5 Main Experiments
We verify the empirical effectiveness of our theoretical analyses. In detail, we demonstrate that controlling the upper bound of the gradient norms also prevents loss and gradient spikes. To assess efficacy in the real situation, we focus on the methods initialized with the widely used method (Shoeybi et al., 2020; Le Scao et al., 2022) in main experiments [^9] The Xavier initialization, which does not satisfy the small sub-layers condition as shown in Table 1, is not widely used for LLMs in recent years. Appendix B shows that the performance of Xavier initialization is worse and fails to avoid spikes..
@@ 5.1 Datasets
We used C4 (Raffel et al., 2020) that consists of clean English texts extracted from Common Crawl [^10] https://commoncrawl.org/ as our LLM pre-training corpus. We also used the separated part of C4 as our validation data. We used GPT-2 vocabulary (Radford et al., 2019) that contains Byte Pair Encoding (BPE) subword units (Sennrich et al., 2016) as our vocabulary. To evaluate each method, we computed perplexity on WikiText (Merity et al., 2017) and LAMBADA (Paperno et al., 2016) datasets.
@@ 5.2 Model Configurations
As described in Section 2, we used the Pre-LN Transformer architecture. We set the number of layers $N=24$, and varied $d$ to adjust the total number of parameters to 350 million (350M) and 1.7 billion (1.7B). We set the learning rate ($\mathrm{lr}$) $5.0\times 10^{-4}$. Section 6.1 shows experiments in varying the learning rate. Appendix A describes more details on the experimental configuration.
(a) 350M parameters.
(b) 1.7B parameters.
Figure 3: Loss curves of each method in validation data.
(a) 350M parameters.
(b) 1.7B parameters.
Figure 4: Gradient norms of each method during the training.
We compared the following methods. We put $\checkmark$ before the method name if the method satisfies both conditions to suppress the upper bound. Moreover, Table 1 summarizes whether each method satisfies each condition.
@@
This is the standard configuration for the LLM pre-training. Since this configuration does not suppress the upper bound of the gradient norms, the loss spike is likely to occur.
@@
where $\gamma$ is a hyper-parameter and $\mathrm{Detach}$ detaches an input from the computational graph. We assign $0.1$ to $\gamma$ as in Zeng et al. (2023). Zeng et al. (2023) indicated that this method empirically prevents the loss spike. However, this method does not satisfy the condition on large shortcut, and thus, we show that this method does not completely solve the loss spike.
@@ 5.3 Results
Table 2 shows the perplexities of each method on WikiText and LAMBADA. This table shows that Embed LN and Scaled Embed achieved comparable performance. This result implies that methods have no significant difference from each other in their performance if each method prevents loss and gradient spikes. In contrast, the perplexities of Vanilla and Embed Detach are worse except for Vanilla with 350M parameters in LAMBADA, and the difference in the performance is larger in a large amount of parameters. This result implies that addressing spikes has a more serious influence on the performance as the parameter size gets larger. We discuss this matter in more detail in Section 6.1.
6 Discussions on Other Configurations
6.1 Varying Learning Rate
Le Scao et al. (2022) reported that the stable method, such as Embed LN, was worse than Vanilla. However, in Section 5, the stable methods, Scaled Embed and Embed LN, achieved better performance than Vanilla in the 1.7B parameter configuration. We suppose that the difference in the learning rate causes this gap in findings. In this section, we tried to train Vanilla and Scaled Embed with larger and smaller learning rates: $\mathrm{lr}=1.0\times 10^{-3}$ and $1.0\times 10^{-4}$ respectively.
(a) 350M with $\mathrm{lr}=1.0\times 10^{-3}$.
(b) 350M with $\mathrm{lr}=5.0\times 10^{-4}$.
(c) 350M with $\mathrm{lr}=1.0\times 10^{-4}$.
(d) 1.7B with $\mathrm{lr}=1.0\times 10^{-3}$.
(e) 1.7B with $\mathrm{lr}=5.0\times 10^{-4}$.
(f) 1.7B with $\mathrm{lr}=1.0\times 10^{-4}$.
@@ 6.1 Varying Learning Rate
Model lr $\mathrm{lr}\ 1.0\times 10^{-3}$ lr $\mathrm{lr}\ 5.0\times 10^{-4}$ lr $\mathrm{lr}\ 1.0\times 10^{-4}$ lr $\mathrm{lr}\ 1.0\times 10^{-3}$ lr $\mathrm{lr}\ 5.0\times 10^{-4}$ lr $\mathrm{lr}\ 1.0\times 10^{-4}$
350M parameters
Vanilla 29.96 30.35 34.51 25.12 24.73 32.49
Scaled Embed 28.09 29.86 35.66 22.03 24.37 37.14
1.7B parameters
Vanilla N/A 22.58 23.54 N/A 15.22 16.17
Scaled Embed 20.95 21.29 23.78 12.26 12.53 15.39
6.2 Varying Sequence Length
As described in Section 2.1, Transformers add the output of each sub-layer to the shortcut. Since the standard deviation of the self-attention layer tends to decrease with the length of an input sequence especially at the early stage [^11] See Appendix H for details., a long sequence tends to keep the standard deviation of the shortcut small. Therefore, the long sequence makes the pre-training of Vanilla more unstable.
We conducted experiments with varying the length of the input sequence $L$ from $128$ to $2048$. To use the same number of tokens to update parameters, we adjusted the batch size. Figure 6 shows loss values of Vanilla with each $L$ configuration in the validation data. This figure shows that spikes occurred only in the large $L$, i.e., $1024$ and $2048$. Moreover, the spikes are more likely to occur at the early stage of the pre-training. Therefore, using a short sequence stabilizes the training at the early stage, as reported in Li et al. (2022).

### v4

@@ 1 Introduction
In this research, we provide theoretical analyses focusing on the loss spike problem during LLM pre-training. We identify the upper bound of the gradient norms for the Transformer-based LLMs through analyses on the spectral norms of the Jacobian matrices for the sub-layers. If the upper bound is large, the gradients may spike suddenly, and we assume that this phenomenon causes the loss spike. Then, we indicate that the upper bound is large in the typical setting, such as the widely used implementation, Megatron-LM (Shoeybi et al., 2020), and thus, the loss spike is likely to occur. In addition, to make the upper bound sufficiently small, we introduce two conditions: (1) initializing the parameters of sub-layers with a small value and (2) making the standard deviation of each embedding close to 111. The former condition can be satisfied by the widely used initialization method for LLMs (Shoeybi et al., 2020; Le Scao et al., 2022; Biderman et al., 2023). On the other hand, the latter condition was satisfied in the original Transformer by scaling embeddings (Vaswani et al., 2017), but such scaling is missing from recent implementations. To sum up, through theoretical analyses, we re-evaluate several previous techniques in terms of the stabilization of LLM pre-training.
@@ 2.1 Pre-LN Transformer
where $\mathrm{LN}$ is the layer normalization function [^1] We discuss the difference from the architecture using Root Mean Square layer normalization (RMSNorm) (Zhang & Sennrich, 2019) instead of LN in Appendix B.4, and the original Transformer architecture, i.e., Post-LN Transformer in Appendix H.. We call the first terms in Equations (1) and (2), i.e., $x$ and $x^{\prime}$, shortcut. In addition, the feed-forward network ($\mathrm{FFN}$) and multi-head self-attention ($\mathrm{Attn}$) are defined as follows [^2] To simplify equations, we omit bias terms.:
@@ 3 Analyses on Gradients of Sub-Layers
Let $x$ and $x^{\prime}$ be the input and intermediate vectors of each layer. Moreover, let $W_{*}$ denote the model parameter matrix in each layer. We assume that $x$, $x^{\prime}$, and $W_{*}$ for all layers follow a normal distribution with a mean of 0, i.e., $\mu=0$.
This assumption is valid when we initialize parameters with the normal distribution, the number of heads in Equation (4) is 111, and $\mathcal{F}$ is an identity function. Empirically, the outputs of each sub-layer are close to the normal distribution as illustrated in Appendix D.
@@ 3.1 Jacobian Matrix of FFN
$\bigg{\|}\frac{\partial y}{\partial x^{\prime}}\bigg{\|}_{2}$ $=\bigg{\|}\frac{\partial(x^{\prime}+\mathrm{FFN}(\mathrm{LN}(x^{\prime})))}{\partial x^{\prime}}\bigg{\|}_{2}=\bigg{\|}I+\frac{\partial(\mathrm{FFN}(\mathrm{LN}(x^{\prime})))}{\partial x^{\prime}}\bigg{\|}_{2}$ (8)
@@
Regarding the $\mathrm{FFN}$ part, we assume that the activation function $\mathcal{F}$ is an identity function [^3] Appendix E discusses the case where we use the $\mathrm{ReLU}$, $\mathrm{SiLU}$, and $\mathrm{SwiGLU}$ as the activation functions, which leads to the same conclusion. to simplify the discussion. Under this assumption, the following equation holds:
$\bigg{\|}\frac{\partial\mathrm{FFN}(\mathrm{LN}(x^{\prime}))}{\partial\mathrm{LN}(x^{\prime})}\bigg{\|}_{2}=\|W_{2}W_{1}\|_{2}.$ (10)
Therefore, we can straightforwardly derive the relation $\|W_{2}W_{1}\|_{2}\leq\|W_{1}\|_{2}\|W_{2}\|_{2}$ from the submultiplicativity of the spectral norm. Furthermore, let $\sigma_{1}$ and $\sigma_{2}$ be the standard deviations of $W_{1}$ and $W_{2}$, respectively. From Assumption 1, the spectral norms of $W_{1}$ and $W_{2}$ are obtained by their standard deviations and dimensions (Vershynin, 2018), i.e., $\|W_{1}\|_{2}\approx\sigma_{1}(\sqrt{d}+\sqrt{d_{\mathrm{ffn}}})$ and $\|W_{2}\|_{2}\approx\sigma_{2}(\sqrt{d}+\sqrt{d_{\mathrm{ffn}}})$. Finally, we can express an upper bound of the spectral norms of the Jacobian matrices for $\mathrm{FFN}$ as the following inequality:
@@
where the right-hand side has the relation $\sigma_{1}\sigma_{2}(\sqrt{d}+\sqrt{d_{\mathrm{ffn}}})^{2}\approx\|W_{1}\|_{2}\|W_{2}\|_{2}$.
@@
We consider the variance ($\mathrm{var}$) of each element in the matrix $zz^{\top}$. Since $z_{i}z_{i}$ follows $\mathcal{X}^{2}$ with 111 degree of freedom, and $z_{i}z_{j}(i\neq j)$ is the multiplication of two independent values following the standard normal distribution, the variances are as follows:
@@ 3.2 Jacobian Matrix of Self-Attention
Let $Z(\cdot)=\mathrm{concat}(\mathrm{head}_{1}(\cdot),...,\mathrm{head}_{h}(\cdot)))$ and let $J^{Z}$ be the Jacobian of the $Z(\cdot)$ [^4] We discuss the detail of $J^{Z}$ in Appendix G., we can rewrite the spectral norm of the Jacobian matrix of Attn as:
@@ 4 Conditions to Avoid Spikes
However, in this situation, the standard deviation of the shortcut is also too small. For example, at shallow layers, the standard deviation is close to $\sqrt{\frac{2}{5d}}$ because the embedding matrix is also initialized by $\mathcal{N}(0,\sigma^{2})$ where $\sigma=\sqrt{\frac{2}{5d}}$. Therefore, to increase the standard deviation of the shortcut, we make the standard deviation of each embedding close to 111666Based on Equations (15) and (20), the upper bound becomes small as the standard deviation of the shortcut increases. However, a too large value degrades the performance empirically as described in Appendix C.. To achieve this, we introduce two kinds of modification: “Scaled Embed” and “Embed LN” [^7] We can satisfy the condition by initializing embeddings with the normal distribution $\mathcal{N}(0,\sigma^{2})$ where $\sigma=1$, but we do not adopt this strategy in this study because we use the same initialization method in our experiments.. The Scaled Embed scales embeddings with an appropriate value. For example, we multiply embeddings by $\sqrt{d}$, which was used in the original Transformer paper (Vaswani et al., 2017) [^8] Although the original Transformer paper introduced this operation, recent implementations ignore this. Therefore, LLMs trained with recent implementations do not satisfy large shortcut., and then the standard deviations of embeddings become $\sqrt{\frac{2}{5}}$. The Embed LN applies the LN to embeddings. In fact, Le Scao et al. (2022) reported that the Embed LN strategy prevents loss spikes empirically. These two methods are presented as verification examples rather than proposed methods, and alternative approaches could be employed if the conditions are met.
@@
To demonstrate the actual values of the upper bound described in Equation (15), we take the model with 1.7 billion parameters as an example. In addition to the widely used initialization for LLMs (Vanilla) and the above two modifications: Scaled Embed and Embed LN, we compare Xavier Init, which initializes all parameters with the Xavier initialization (Glorot & Bengio, 2010), as the situation where we do not scale $W_{2}$ and $W_{O}$ based on the number of layers. Figure 2 shows the values of Equation (15) for each layer at the beginning of the pre-training. This figure indicates that the methods without suppressing the upper bound, i.e., Xavier Init and Vanilla, rapidly increase the values especially in shallow layers. In contrast, Scaled Embed and Embed LN keep small values. In summary, to make the upper bound of the gradient norms small for the stabilization of the LLM pre-training, we have to satisfy two conditions: (1) small sub-layers; initializing the parameters of sub-layers with a small value and (2) large shortcut; making the standard deviation of each embedding close to 111.
5 Experiments
We verify the empirical effectiveness of our theoretical analyses. In detail, we demonstrate that controlling the upper bound of the gradient norms also prevents loss and gradient spikes. To assess efficacy in the real situation, we focus on the methods initialized with the widely used method (Shoeybi et al., 2020; Le Scao et al., 2022) in main experiments [^9] The Xavier initialization, which does not satisfy the small sub-layers condition, is not widely used for LLMs in recent years. Appendix B.1 shows that the performance of Xavier initialization is worse and fails to avoid spikes..
@@ 5.1 Datasets
We used C4 (Raffel et al., 2020) that consists of clean English texts extracted from Common Crawl as our LLM pre-training corpus. We also used the separated part of C4 as our validation data. We used GPT-2 vocabulary (Radford et al., 2019) that contains Byte Pair Encoding (BPE) subword units (Sennrich et al., 2016) as our vocabulary. To evaluate each method, we computed perplexity on WikiText (Merity et al., 2017) and LAMBADA (Paperno et al., 2016) datasets.
@@ 5.2 Model Configurations
As described in Section 2, we used the Pre-LN Transformer architecture. We conduct experiments with two parameter sizes: 350 million (350M) and 1.7 billion (1.7B). We set the learning rate ($\mathrm{lr}$) $5.0\times 10^{-4}$. Appendix A describes more details on the experimental configuration. We compared the following methods. We put $\checkmark$ before the method name if the method satisfies both conditions to suppress the upper bound.
@@
This is the standard configuration for LLM pre-training. Since this configuration does not suppress the upper bound of the gradient norms, the loss spike is likely to occur.
@@
where $\gamma$ is a hyper-parameter and $\mathrm{Detach}$ detaches an input from the computational graph. We assign 0.10.10.1 to $\gamma$ as in Zeng et al. (2023). Zeng et al. (2023) indicated that this method empirically prevents the loss spike. However, this method does not satisfy the condition on large shortcut, and thus, we show that this method does not completely solve the loss spike.
@@ 5.3 Main Results
Table 1: Perplexities of each method.
Table 1 shows the perplexities of each method on WikiText and LAMBADA. This table shows that Embed LN and Scaled Embed achieved comparable performance. This result implies that methods have no significant difference from each other in their performance if each method prevents loss and gradient spikes. In contrast, the perplexities of Vanilla and Embed Detach are worse except for Vanilla with 350M parameters in LAMBADA, and the difference in the performance is larger in a large amount of parameters. This result implies that addressing spikes has a more serious influence on the performance as the parameter size gets larger. We conduct experiments on a larger model in the following subsection. Moreover, we discuss other settings in Appendix B.
(a) 350M parameters.
(b) 1.7B parameters.
Figure 3: Loss curves of each method in validation data.
(a) 350M parameters.
(b) 1.7B parameters.
Figure 4: Gradient norms of each method during the training.
5.4 Results on 13B Parameter Model
We conduct experiments on pre-trainings of 13B parameter models to indicate that our provided conditions can stabilize a model with many more parameters than the ones discussed in Section 5.3. Due to the limitation of our computational budget, we focused on the comparison between Vanilla and Scaled Embed. We tried two learning rates: $3.0\times 10^{-4}$ and $1.0\times 10^{-4}$. Appendix A describes more details about hyper-parameters.
Figure 5 shows the loss values of each configuration in validation data. As shown in (a) of this figure, the loss value of Vanilla rose from approximately 10000 steps in $\mathrm{lr}=3.0\times 10^{-4}$. Then, the gradient of this model became too large to continue its pre-training. In contrast, the loss value of Scaled Embed consistently decreased. This result indicates that Scaled Embed stabilized the pre-training of the model with a large number of parameters.
(a) $\mathrm{lr}=3.0\times 10^{-4}$
(b) $\mathrm{lr}=1.0\times 10^{-4}$.
@@ 5.4 Results on 13B Parameter Model
Vanilla N/A 15.12 N/A 6.50
Scaled Embed 14.47 15.25 5.97 6.53
Model PIQA $\uparrow$ OpenBookQA $\uparrow$ HellaSwag $\uparrow$ WinoGrande $\uparrow$ ARC-easy $\uparrow$ ARC-challenge $\uparrow$
$\mathrm{lr}=3.0\times 10^{-4}$
Vanilla N/A
Scaled Embed 78.94 39.20 71.03 63.77 60.31 35.49
$\mathrm{lr}=1.0\times 10^{-4}$
Vanilla 77.80 38.40 69.10 61.64 57.95 33.53
Scaled Embed 78.07 39.20 68.92 62.59 58.92 33.19
Table 3: Performance of each method with 13B parameters on standard tasks.
Table 2 shows the perplexities of each configuration in evaluation data. This table indicates that we can achieve better performance when we use a larger learning rate. In addition, the perplexities of Scaled Embed were comparable to ones of Vanilla when we used the small learning rate: $\mathrm{lr}=1.0\times 10^{-4}$. Table 3 shows the performance of each configuration on the standard benchmark datasets: PIQA (Bisk et al., 2020), OpenBookQA (Mihaylov et al., 2018), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021) and ARC easy and challenge (Clark et al., 2018). This table shows results consistent with the evaluation in terms of perplexity. These results imply that our provided conditions have no considerable risk in pre-training. Thus, we have to satisfy large shortcut in addition to small sub-layers to stabilize the pre-trainings of LLMs.